Tohban Report 2016-03-23
Jump to navigation
Jump to search
Tohban Reports | |
---|---|
Start Date: | 16 Mar 2016 |
End Date: | 23 Mar 2016 |
Tohban: | Nicole Duncan |
Tohban email: | nicoleduncan@ssl.berkeley.edu |
Next Tohban: | Martin Fivian |
List all reports |
Solar Activity
Solar activity this week was very low.
How many GOES flares occurred? Flares above B, C, M, X class were 8 3 0 0
RHESSI flare list contains 0 over the time range 16-Mar-16 22Mar-16
Memory Management
SSR filled up this week to ~80% due to ADP testing. Over the past 3 days, it is essentially hovering around the 60 - 80% data range. Martin will look into adjusting the ADC tables.
Spacecraft Status
The spacecraft is currently in the middle of an anneal procedure. Cool down going well, with the cold tip falling at .3 deg /orbit. Low accelerations around 12mg. We decided to progress with a narrower ~10% margin on the guidelines. This was done on Friday and the cryocooler is responding fine.
Data Gaps
N/A
Detector issues
The detectors are currently off.
Spacecraft Management
Decimation | N/A |
HLAT Decimation | N/A |
Night time data (fronts) | N/A |
Night time data (rears) | N/A |
Require extra passes? | N/A |
Requirement for moving pointer? | N/A |
Attenuator operation | N/A |
Detector problems? | N/A |